OpenStudy (anonymous):

11mL of 1 mol/L NaCl and 8mL of 0.10 mol/L AgNO3 they react. Find theoretical yield.

OpenStudy (anonymous):

Can you figure out the balanced equation?

OpenStudy (anonymous):

i did NaCl = AgNO3 -> NaNO3 + AgCl

OpenStudy (anonymous):

NaCl + *

OpenStudy (anonymous):

that's balanced^

OpenStudy (anonymous):

alright now you have to find the limiting reagent can you do that?

OpenStudy (anonymous):

i did i got AgNO3 as the limiting

OpenStudy (anonymous):

now how do i get the theoretical yield if the mole ratio is 1:1

OpenStudy (anonymous):

Does it tell you what product to take it to? if not pick one

OpenStudy (anonymous):

no its one of those questions where you just have to find theoretical yield :O

OpenStudy (anonymous):

\[8 ml AgNO_3(\frac{0.10 mole AgNO_3}{1000 ml AgNO_3})(\frac{1 mole AgCl}{1 mole AgNO_3})(\frac{MM AgCl}{1 mole AgCl})=theoridcal yeild\] where i put in MM put the molar mass of AGCL and follow the steps and this should tell you the theoridical yeild